OpenStudy (anonymous):

Can someone explain me this, plz! Solve the system by substitution. x+7y=-2 y=-3x+34 Thx!!

OpenStudy (anonymous):

plug y into x, so we have x+7(-3x+34)=-2 now solve for x x-21x+238=-2 -20x=-240 x=12 plug x back into one of the equation to get y

OpenStudy (anonymous):

x+7y=-2 y=-3x+34 -- x=-2-7y * -- * y=-3(-2-7y)+34 -- * y=6+21y+34 -20y=40 y=-2 and x+7(-2)=-2 x-14=-2 x=12 *** x=12 y=-2

OpenStudy (anonymous):

resolves?

OpenStudy (anonymous):

Thx!!